Operations Report – April 7th

Operations Report 7-APR-2019

SOL: n/a

Name of person filing report: Shannon Rupert

Non-nominal systems: generator

Notes on non-nominal systems: generator needs maintenance and replacement of alternator. I plan to change the oil and replace the alternator tomorrow. That will leave only the coolant flush and hose replacement, which I don’t quite know how to do.

Generator (hours run): Off at 8 am, not sure when it was turned on last evening

Solar— 85 SOC % (Before generator is run at night). Turned generator on at 8 pm. Neither charger nor inverter shows as operational. That is consistent with my gradual loss of power throughout the day. I should have seen 100% until late in the afternoon, instead we were down to 95 by noon. So I will investigate this further in the morning.

Diesel Reading – 48 %

Propane Reading – 50 %

Ethanol Free Gasoline – 0 gallons.

Water (auxillary tank) – 0 gallons

Water (static tank) – 75 gallons

Auxillary to Static tank transfer– no

Gallons transferred: No

Water in GreenHab – unknown gallons

Water (loft) – Static to Loft Pump used – no

Water Meter: not read

Toilet tank emptied: no, It should be done, but there is not enough water. In addition, there is no digester that I can find in the Hab. I will pick up some tomorrow.

Deimos rover used: Still non-operational


Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: ASSIGNED TO DIRECTOR, yes

Hours: 50.7

Beginning charge: 100

Ending charge: 85

Currently charging: yes

Spirit rover used: No

Hours: 86.5

Beginning charge: 100

Ending charge: 100

Currently charging: yes, but parked in the wrong place

Opportunity rover used: No

Hours: 56.4

Beginning charge: 100

Ending charge: 100

Currently charging: yes

Curiosity rover used: No

Hours: 88.8

Beginning charge: 100

Ending charge: 100

Currently charging: yes, but parked in the wrong place

Notes on rovers: All four rovers require cleaning. Crew 208 reported that the mud was there when they arrived. I will work on getting them cleaned up this week. They should have been cleaned during checkout. Sojourner is also very muddy.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3)

Reason for use: None were used.

Oil Added? No

ATV Fuel Used: 0 Gals. A couple of gas cans have a small amount of gas in them but I don’t know if it is ethanol free so I will just get more gas

# Hours the ATVs were Used today:0

Notes on ATVs: Two of the ATVs (Honda and 350.2) are absolutely filthy. Two (300 and 350.1) are fine. 350.3 is parked next to Sojourner. Is something wrong with it?

HabCar used and why, where? No, although I will need to get water with it tomorrow. It is also very muddy. There is also no shovel or jumper cables in it.

CrewCar used and why, where? The crew took it back to Grand Junction today. It was very dirty with clumped mud. David got most of it off but it needs to be given a good cleaning. Julielynn noted that when they originally picked it up it was not filled with gas. We should remind crews to read instructions in the Handbook before using the car.

General notes and comments: This report contains my assessment of the current state of the campus. It is not meant to criticize anyone, just to let everyone know my judgement on the state of systems.

Overall the campus is in very good shape. Some things have been neglected, and will need attention. There is, as usual, a collection of strange clutter and things out of place. My favorite is the four by four post underneath the stairs with some other junk. Like, why?

The damaged cattle panels and the rock pile remain from the beginning of the season. I will be able to move the rock pile, but will need help with the panels.

There are two pumps on the ground by my house. Are they working or broken?

It was a lovely day so I opened all the doors and aired all the buildings. We should have one more nice calm warm day before it gets cold and windy again.

The main road (Cow Dung) is in really poor condition and is in worse shape than when I left in January. Only raised vehicles can safely travel it and all rover travel should be done in MAX mode until it is improved. I’ve had no luck getting the county to maintain it. I will contact the BLM for assistance in this request. I also will suggest we ask the county again for maintenance before the URC or people won’t be able to get here.

Summary of internet: My internet had 0 bytes remaining for the next two weeks. 10 GB were added this afternoon. Thanks, Bernard.

Summary of suits and radios: Nothing done today

Summary of Hab operations: The upper deck is ok. It needs a complete cleaning. The kitchen cabinets are dirty and there is still junk on the kitchen shelves (Atila!). The lower deck is dirty under the stairs, like no one has cleaned it in several crews. There are no jumper cables in the EVA room and it is dirty. There are two batteries in the EVA room and I don’t know what they are for and why they are in there.

Summary of GreenHab operations: Both the heater and the fan were on this morning. I turned the fan off and the heater down to 10 degrees C. The GreenHab is neat and tidy, but there are a lot of pots from crew projects and failed crops that have not been taken care of. I will take care of this in the next few days. The GreenHab
Coordinator, David Murray, will be here next week and he can decide what and if we should continue planting for the rest of the season. The germination tent is also in the GreenHab and was not returned to the ScienceDome after crew use as it should have been. It’s light but awkward and I will need help to move it.

Summary of ScienceDome operations: The Science Dome is in relatively good shape. One table is missing and there is trash that needs to be removed. General organization and clean up must be done. The floor is not clean. There are some unknown samples in biohazard bins out by the generator and I don’t know what they are.

Summary of RAM operations: Cluttered and messy. I will remove the extra materials and clean it up. On the south side of the building, the paint is peeling off above and on the side near the roll up door. My guess is this is due to the ice build up and then melt we had so much this winter

Summary of any observatory issues: Nothing done today.

Summary of health and safety issues: Nothing to report

Questions, concerns and requests to Mission Support: Please confirm that this report has been posted on the website. It is important that we retain these data.

Copyright © The Mars Society. All rights reserved. | Main Site